# All-in-one: Understanding and Generation in Multimodal Reasoning with the MAIA Benchmark

### v2

@@ B.2 Task1: Visual Statement Verification
Black video InternVL2 0.68 0.64 0.88 0.89 0.69 0.61 0.96 0.52 0.59 0.54 0.55 0.67 0.60
Llava-Next-Video 0.50 0.49 0.52 0.48 0.51 0.52 0.45 0.50 0.51 0.51 0.49 0.51 0.48
Llava-oneVision 0.59 0.55 0.48 0.92 0.61 0.38 0.97 0.52 0.53 0.50 0.52 0.60 0.51
Qwen-2.5-VL 0.73 0.68 0.86 0.87 0.75 0.72 0.98 0.56 0.62 0.61 0.65 0.73 0.70
1-Frame InternVL2 0.75 0.80 0.87 0.69 0.71 0.77 0.89 0.65 0.80 0.65 0.82 0.67 0.69
Llava-Next-Video 0.61 0.68 0.76 0.49 0.63 0.75 0.40 0.59 0.65 0.59 0.70 0.56 0.53
Llava-oneVision 0.76 0.79 0.78 0.89 0.74 0.76 0.91 0.68 0.81 0.64 0.81 0.70 0.67
Qwen-2.5-VL 0.79 0.81 0.81 0.91 0.75 0.82 0.70 0.82 0.96 0.67 0.82 0.69 0.73
32-Frames InternVL2 0.79 0.83 0.85 0.77 0.75 0.84 0.84 0.75 0.83 0.69 0.86 0.66 0.76
Llava-Next-Video 0.52 0.56 0.57 0.42 0.57 0.61 0.32 0.52 0.59 0.52 0.56 0.51 0.48
Llava-oneVision 0.81 0.87 0.78 0.88 0.76 0.88 0.85 0.80 0.85 0.71 0.87 0.65 0.80
Qwen-2.5-VL 0.84 0.89 0.80 0.89 0.78 0.86 0.92 0.82 0.88 0.83 0.90 0.75 0.81
@@
Table 11: Visual Statement Verification (Task $1$): accuracy of correct choices across reasoning categories.
Models Total Agreement Majority Agreeement
TS FS TS FS
Black video InternVL 17.83% 0.58% 63.79% 17.80%
LLaVa-neXT-Video 0.21% 0.29% 63.20% 36.30%
LLava-oneVision 13.91% 1.71% 57.20% 27.18%
Qwen2.5-VL 22.25% 0.54% 66.25% 10.96%
1-Frame InternVL 27.58% 0.70% 60.42% 11.30%
LLaVa-neXT-Video 8.30% 0.96% 68.70% 22.04%
LLava-oneVision 31.55% 0.71% 56.62% 11.12%
Qwen2.5-VL 35.92% 0.25% 55.33% 8.50%
32-Frame InternVL 31.25 % 0.30% 60.04% 8.41%
LLaVa-neXT-Video 2.79% 1.46% 62.25% 33.5%
LLava-oneVision 37.71% 0.21% 54.33% 7.75%
Qwen2.5-VL 43.62% 0.21% 50.92% 5.25%
Table 12: Task 1: consistency of TS and FS across the 8-answer pool. Total agreement is 8/8, majority is between 4 and 7. There are 00 cases with agreement < 4.
We further examined the accuracy variability across the 8 pools, with Figure 17 providing a detailed examination of trends within each category. This analysis revealed notable fluctuations in accuracy for certain categories such as Uncertainty, Spatial, and Temporal categories, pointing to a lack of stability and highlighting the complexity of tasks . For instance, some models, like LLaVA-NeXT-Video, exhibited more significant variability, showcasing their challenge in handling these categories. Despite these fluctuations, Qwen2.5-VL emerged as the top performer overall, maintaining the most consistent accuracy across pools. Similar conclusions can be reached through visual inspection of Figures 15 and 16.
Figure 17: Performance trends of the four models. Each plot shows the accuracy trend as a function of the 8 different pools.
B.3 Task2
We employed two evaluation approaches for the generation task: similarity-based metrics and the LLM-as-a-judge approach. The latter proved to be the most suitable for our purposes, given its flexibility and ability to handle open-ended responses effectively. The use of 8 reference answers was instrumental in this process. Inspired by the excellent work of Lee et al. (2024a) and Mañas et al. (2024), we leveraged these references to assess model-generated responses without enforcing strict syntactic similarity. This approach allows for a more nuanced evaluation, prioritizing semantic alignment over surface-level structure. Notably, these reference answers are slightly different but all correct, capturing a range of valid perspectives. This ensures a more robust and reliable evaluation, as it accounts for the natural variability in human responses and the multiple ways a model might correctly express the same information. To ensure further robustness, we adopted GPT-4o as our evaluation model, prompting it to determine correctness by checking whether a candidate answer aligns with at least one reference. Most importantly, we only proposed this method after verifying its high correlation with two human annotators (Fleiss’ Kappa: 0.82), ensuring its reliability. For transparency, Figure 18 represent the prompt used for the automatic evaluation of VLMs’ answers. Figure 19 represetns models’ fingerprint through MAIA’s reasoning categories, specifically illustrating models’ performance in the NLG Task 2.
Explicit Implicit Partial Total Partial Total Duration Partial
Black video InternVL2 0.37 0.41 0.43 0.60 0.30 0.68 0.43 0.08 0.21 0.23 0.36 0.25 0.55 0.25
Llava-Next-Video 0.27 0.40 0.47 0.47 0.30 0.40 0.33 0.51 0.21 0.12 0.16 0.04 0.12 0.16
Llava-oneVision 0.40 0.60 0.73 0.68 0.29 0.60 0.60 0.28 0.26 0.24 0.36 0.18 0.37 0.23
Qwen-2.5-VL 0.35 0.35 0.37 0.69 0.08 0.54 0.74 0.23 0.24 0.19 0.30 0.19 0.41 0.20
1-Frame InternVL2 0.44 0.48 0.67 0.65 0.21 0.65 0.60 0.10 0.33 0.38 0.39 0.47 0.53 0.35
Llava-Next-Video 0.32 0.24 0.36 0.56 0.20 0.46 0.60 0.29 0.18 0.32 0.24 0.27 0.20 0.19
Llava-oneVision 0.50 0.56 0.62 0.78 0.15 0.66 0.74 0.37 0.39 0.54 0.39 0.51 0.54 0.33
Qwen-2.5-VL 0.51 0.50 0.62 0.76 0.32 0.60 0.79 0.40 0.35 0.50 0.38 0.53 0.55 0.38
32-Frames InternVL2 0.49 0.50 0.58 0.68 0.28 0.64 0.62 0.11 0.45 0.48 0.47 0.51 0.57 0.46
Llava-Next-Video 0.33 0.33 0.42 0.38 0.16 0.42 0.48 0.27 0.24 0.37 0.29 0.39 0.32 0.29
Llava-oneVision 0.53 0.63 0.71 0.79 0.11 0.65 0.79 0.23 0.55 0.56 0.51 0.62 0.40 0.46
Qwen-2.5-VL 0.61 0.68 0.75 0.80 0.43 0.60 0.85 0.55 0.55 0.60 0.50 0.70 0.54 0.53
Table 13: Open-ended Visual Question Answering (Task $2$): accuracy of correct answers with LLM-as-a-judge with Causal sub-categories specification.
@@ B.3 Task2
B.4 Case study: Multimodal Hallucinations
As part of a preliminary error analysis focusing on the Out-of-Scope category, we observed that a significant portion of the errors made by the model could be attributed to multimodal hallucinations. Particularly interesting was the discovery of counterintuitive clashes between the two tasks in our benchmark. In several instances, the model successfully solved Task 1 (i.e., True Statement Selection), in some cases achieving also full consistency within the pools (e.g., 8/8 correct selections), yet failed Task 2 (i.e., open-ended NLG), generating hallucinated content. For example, given a video in which a young girl is frightened by an insect in her home and seeks her mother’s help to remove it, our best-performing model (i.e., Qwen2.5VL) consistently selected the correct true statement in all eight pairs [^13] True Statements are the first in each pair:

### v3

@@ B.2 Tasks Details
Black video InternVL2 8B 0.68 0.64 0.88 0.89 0.69 0.61 0.96 0.52 0.59 0.54 0.55 0.67 0.60
InternVL3 78B 0.77 0.81 0.88 0.60 0.80 0.84 0.99 0.64 0.71 0.68 0.71 0.78 0.82
Llava-Next-Video 7B 0.50 0.49 0.52 0.48 0.51 0.52 0.45 0.50 0.51 0.51 0.49 0.51 0.48
Llava-oneVision 7B 0.59 0.54 0.48 0.92 0.61 0.38 0.97 0.52 0.53 0.50 0.52 0.60 0.51
Qwen-2.5-VL 7B 0.73 0.68 0.86 0.87 0.75 0.72 0.98 0.56 0.62 0.61 0.65 0.73 0.70
Qwen-2.5-VL 72B 0.69 0.71 0.78 0.95 0.75 0.57 0.99 0.53 0.62 0.54 0.57 0.67 0.63
1-Frame InternVL2 8B 0.75 0.80 0.87 0.69 0.71 0.77 0.89 0.65 0.80 0.65 0.82 0.67 0.69
InternVL3 78B 0.81 0.86 0.88 0.81 0.83 0.84 0.93 0.74 0.87 0.68 0.84 0.70 0.77
Llava-Next-Video 7B 0.61 0.68 0.76 0.49 0.63 0.75 0.40 0.59 0.65 0.59 0.70 0.56 0.53
Llava-oneVision 7B 0.76 0.79 0.78 0.89 0.74 0.76 0.91 0.68 0.81 0.64 0.81 0.70 0.67
Qwen-2.5-VL 7B 0.79 0.81 0.81 0.91 0.75 0.82 0.70 0.82 0.96 0.67 0.82 0.69 0.73
Qwen-2.5-VL 72B 0.79 0.81 0.81 0.92 0.79 0.74 0.97 0.73 0.84 0.67 0.85 0.70 0.70
32-Frames InternVL2 8B 0.79 0.83 0.85 0.77 0.75 0.84 0.84 0.75 0.83 0.69 0.86 0.66 0.76
InternVL3 78B 0.64 0.81 0.87 0.55 0.71 0.59 0.62 0.61 0.76 0.67 0.71 0.79 0.82
Llava-Next-Video 7B 0.52 0.56 0.57 0.42 0.57 0.61 0.32 0.52 0.59 0.52 0.56 0.51 0.48
Llava-oneVision 7B 0.81 0.87 0.78 0.88 0.76 0.88 0.85 0.80 0.85 0.71 0.87 0.65 0.80
Qwen-2.5-VL 7B 0.84 0.89 0.80 0.89 0.78 0.86 0.92 0.82 0.88 0.83 0.90 0.75 0.81
Qwen-2.5-VL 72B 0.88 0.93 0.86 0.95 0.85 0.89 0.95 0.88 0.92 0.80 0.92 0.69 0.84
@@
Table 5: Visual statement verification (Task 11): accuracy of correct choices across reasoning categories (without aggregation).
Figure 6: Task 1
Figure 7: Task 1 pool-based
@@
B.3 Case study: Multimodal Hallucinations
As part of a preliminary error analysis focusing on the Out-of-Scope category, we observed that a significant portion of the errors made by the model could be attributed to multimodal hallucinations. Particularly interesting was the discovery of counterintuitive clashes between the two tasks in our benchmark. In several instances, the model successfully solved Task 1 (i.e., True Statement Selection), in some cases achieving also full consistency within the pools (e.g., 8/8 correct selections), yet failed Task 2 (i.e., open-ended NLG), generating hallucinated content. For example, given a video in which a young girl is frightened by an insect in her home and seeks her mother’s help to remove it, one of our best-performing model (i.e., Qwen2.5VL) consistently selected the correct true statement in all eight pairs [^16] True Statements are the first in each pair: